OpenStudy (anonymous):

y'' + 4y = {3t, 0 (<=) t (<=)7}, {21, t>7} for unit step equations you do the laplace transform to both sides, but I can't get the right side unit step function to work properly. What I get is: (u(t) - u(t-7) )3t + 21(u(t-7)) and then you apply the corresponding laplace transform help please?

OpenStudy (anonymous):

3t, 0 less than or equal to (t) less than or equal to 7 <- this for piece wise function on the right

OpenStudy (amistre64):

\[L\{y''\}=s^2Y(s)-sy(0)-y'(0)\] right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and Laplace transform of 4y is just 4Y

OpenStudy (amistre64):

\[4L\{y(t)\}=4Y(s)\]

OpenStudy (amistre64):

then the right side ..... we split up and it messes up our pretty laplace tables :)

OpenStudy (anonymous):

i mean i think i set up the unit step functions possibly correct: u(t) -u(t-7) for the first bounds

OpenStudy (anonymous):

then u(t-7) for the second bound?

OpenStudy (amistre64):

\[\int e^{-st}t\ dt=e^{-st}+\frac{1}{s}\int e^{-st}dt\] \[\int e^{-st}t\ dt=e^{-st}+\frac{e^{-st}}{s^2};[0,21]\] \[e^{-21s}+\frac{e^{-21s}}{s^2}-1-\frac{1}{s^2}\] im not sure what a unit step function is tho

OpenStudy (amistre64):

and change 21 to 7 :) this old man cant see too well

OpenStudy (anonymous):

hmm but we still havent taken into account that the function = 3t at values of t [0,7]

OpenStudy (amistre64):

i forgot a step in my integration by parts; or at least mixed something up on the beginning and i forgot the 3 .... ugh :/

OpenStudy (amistre64):

redo lol

OpenStudy (anonymous):

Seems like I am 19 minutes to late.

OpenStudy (anonymous):

take your time :D

OpenStudy (anonymous):

not sure how to summarize unit step functions, but there a way to characterize the piecewise function without all the inequalities

OpenStudy (amistre64):

\[\int 3t\ e^{-st}\ dt\to\ 3\frac{e^{-st}}{-s}+\frac{1}{s}\int e^{-st}dt\] \[3\frac{e^{-st}}{-s}+\frac{1}{s}(\frac{1}{-s} e^{-st})\] \[3\frac{e^{-st}}{-s}-\frac{e^{-st}}{s^2}\] oh theres plenty of room in here

OpenStudy (amistre64):

i think that 3 spose to distrribute thru ....

OpenStudy (anonymous):

isnt the fact that its 3t only on the interval [0,7] supposed to shift the function over by seven or something?

OpenStudy (amistre64):

dunno, havent gotten to shifts yet in class. so far when weve done these we just integrate it on the short interval

OpenStudy (anonymous):

hmm i'll try what you have there

OpenStudy (amistre64):

when i do it on paper and dont have to focus on the latex coding ... i get: \[L\{3t\};[0,7]\to\ \frac{-21se^{-7s}-3e^{-7s}+3}{s^2}\]

OpenStudy (amistre64):

\[L\{21\}=\frac{21e^{-7s}}{s}\]

OpenStudy (anonymous):

how exactly did you do that? O_o